# COMPASS: Cooperative Multi-Agent Persistent Monitoring using Spatio-Temporal Attention Network

### v1

@@ I Introduction
Persistent monitoring requires a team of mobile agents to visit spatially dispersed, dynamically evolving targets often enough to keep an accurate belief over their states. Representative missions include tracking marine mammals, mapping algal blooms, inspecting crop stress, and assessing post-disaster damage, where data must be gathered under tight limits on energy, communication, and sensing range [1, 2]. A practical algorithm therefore has to balance two coupled objectives: minimising the time between consecutive visits to each location, thus guaranteeing spatial coverage, while also selecting viewpoints that best reduce predictive uncertainty as targets move over time, thus ensuring temporal tracking.
@@
Figure 1: Illustration of the COMPASS Framework in a Simulated Multi-Agent Wildlife Monitoring Task. (Left) A composite view illustrating the planner’s output, where agents’ belief maps of target locations are overlaid with their historical trajectories. The heat map indicates areas of high uncertainty that the agents are incentivized to explore. (Right) Deployment within AirSim-based 3D simulation. Multiple UAVs cooperatively monitor dynamically moving animal targets, each modeled with unique 3D models.
Earlier research illuminates separate facets of this challenge but stops short of a unified solution. STAMP leverages temporal attention to trade exploration for revisit frequency but assumes a single robot [3]. MA-G-PPO extends attention mechanisms to multi-agent reinforcement learning (MARL) yet depends on continuous communication and a coarse global occupancy grid [4]. GALOPP accounts for sensing, localisation, and connectivity constraints in heterogeneous teams, though it lacks a principled way of incorporating historical observations when choosing future actions [5]. By jointly modelling uncertainty with GPs and coupling spatial and temporal reasoning inside one transformer, COMPASS closes this gap and produces decentralised policies that adapt online to moving targets.
@@ III Problem Setup
We consider the task of multi-agent persistent monitoring, where a team of $M$ autonomous agents monitors $N$ mobile targets $\mathcal{T}=\{T_{1},...,T_{N}\}$ within a bounded 2D space $\mathcal{W}\subseteq\mathbb{R}^{2}$. Each target $j$ has an evolving state (e.g. location, field distribution) $x_{j}(t)\in\mathcal{X}_{j}$, governed by unknown or stochastic dynamics. The objective is to maintain accurate belief estimates over all target states throughout the mission.
@@ III-A Belief Modeling with Gaussian Processes
In our implementation, we employ the Matérn kernel [25] with separate spatial and temporal length scales. Sensor observations, represented as binary detections indicating the presence or absence of each target within the sensor’s field of view, are incorporated as data points $((p_{\text{obs}},t_{\text{obs}}),y_{\text{obs}})$ to incrementally update the GPs.
@@ III-C Sensing and Constraints
Figure 2: Our proposed attention-based spatio-temporal neural network for multi-agent continuous monitoring. Agent observations are used to update Gaussian Process (GP) beliefs over time. Historical node features, composed of GP predictions and agent presence information, undergo temporal pooling and initial encoding (Target Encoder). These features are sequentially processed by a Temporal Decoder and a Spatial Encoder, which incorporate masks and positional encodings (PE) in temporal and spatial dimensions, along with current presence embeddings. We fuse distance information before extracting agent-specific features (Feature Extraction). A Pointer Network head (Actor) scores each agent’s neighboring nodes and produces action logits, while a centralized MLP head (Critic) concatenates all agent-specific features and uses an MLP to estimate the shared state-value.
@@ IV-A Problem Modeling and Belief Estimation
At each decision step, agents rely on a belief representation constructed using Gaussian Processes (GPs), as described in Section III-A. These GPs provide probabilistic estimates of each target’s presence across space and time, and we evaluate their posterior mean and variance at every node for both the current time $t$ and a short-term future horizon $t+\delta$. These predictions are concatenated into node-level feature vectors that represent the current and anticipated information landscape. To reflect the deployment state, a binary node presence indicator is appended to each node feature vector, indicating whether it is currently visited by any agent. This representation supports decision-making under uncertainty by informing agents of each other’s locations, enabling coordinated spatio-temporal task allocation via the attention-based policy network.
@@ IV-C Network Architecture
Temporal Decoder
We implement the temporal module as a Transformer decoder. At each decision step, we use the latest timestep feature as the query and attend over the pooled historical features as the memory. To support variable-length histories, we apply an attention mask. We incorporate temporal scale information by adding learnable timestep encodings to the input features. This process yields a temporally contextualized embedding $e_{\text{temp}}(v,t)$ for each node.
@@
To capture spatial correlations, the temporal embeddings are further processed using a Transformer encoder that attends across all nodes in the graph. Each node representation is augmented with spatial information including a positional encoding derived from the eigenvectors of the graph Laplacian and a current presence embedding. The resulting spatial embedding $e_{\text{spat}}(v,t)$ is concatenated with the shortest path distance from each node to the closest agent. A fusion layer then projects this combined representation back to the embedding dimension, resulting in the final node feature $e_{\text{final}}(v,t)$.
@@
For each agent, its current location and the set of $K$ neighboring nodes define its candidate action set. The agent queries the features of these nodes using a single-head attention mechanism. The resulting attention scores are converted into log-probabilities over available actions, forming the policy output $\pi(a|s)$. In parallel, a decentralized value head estimates the state value $V(s)$ based on the agent’s current node embedding. Both the policy and value outputs are derived from the same shared backbone, enabling joint optimization while allowing for decentralized execution at inference time.
@@ IV-D Training Settings
We train our COMPASS policy using Proximal Policy Optimization (PPO), leveraging parallel rollout workers and shared network parameters for actor and critic heads. In each training iteration, $N_{\text{env}}=16$ environment instances are simulated concurrently, each running a full team of agents. Rollouts of fixed length $T=100$ are generated and stored in a centralized experience buffer, with all agent trajectories contributing to joint gradient updates.
The policy is trained using a clipped surrogate objective to ensure stable updates. Let $r_{k}(\theta)$ denote the importance sampling ratio between the current and old policies, and $\hat{A}(k)$ the estimated advantage, the PPO loss is defined as:
@@
where $\epsilon=0.2$. Generalized Advantage Estimation is used with discount factor $\gamma=0.99$ and smoothing parameter $\lambda=0.95$. The value function is trained with mean squared error, and an entropy bonus is added to encourage exploration.
We use the Adam optimizer with an initial learning rate of $10^{-4}$, exponentially decayed by a factor of 0.96 every 64 iterations. Each PPO update runs for 4 epochs per batch. Training proceeds for 50000 environment steps and takes approximately 40 GPU-hours on a cluster equipped with 2 NVIDIA A100 GPUs.
@@ V-A Experimental Setup
We conduct all experiments in simulated environments. Evaluations including those in Table I and Table II, are performed in a scripted simulator with visualization. We also validate our framework in AirSim for high-fidelity simulation, as shown in Fig. 3, which captures realistic spatial dynamics and multi-agent coordination in 3D settings. The workspace $\mathcal{W}$ is a continuous 2D area discretized into a graph $\mathcal{G}$, where $K$ nodes are uniformly sampled in the unit square $[0,1]^{2}$. Each node is connected to its $k=10$ nearest neighbors to ensure connectivity. Unless otherwise specified, we set $K=200$, $r_{\mathrm{sense}}=0.1$ and use a default speed factor of $0.6$, indicating that targets move at $60\%$ of the agents’ speed.
@@ V-B Performance Analysis and Scalability
Figure 3: Evolution of Average Uncertainty (Avg Unc) over Mission Time (seconds) for COMPASS and baseline methods in the default simulation setting ($K=200,M=3,N=8$, Speed Factor 1.0×1.0$\times$1.0 ×). Solid lines represent the mean, shaded areas represent the standard deviation over 20 runs.

### v2

@@ I Introduction
Persistent monitoring requires a team of mobile agents to frequently visit dynamic targets to maintain an accurate, shared belief of their states. This problem is central to applications like environmental sensing and disaster response, which operate under strict resource constraints [1, 2]. A successful strategy must balance spatial coverage—minimizing revisit times—with temporal tracking, selecting viewpoints that best reduce uncertainty as targets move.
@@
Figure 1: Illustration of the COMPASS framework in a simulated multi-agent wildlife monitoring task. (Left) Visualization of three UAV agents’ trajectories and corresponding belief maps over dynamic targets; brighter areas denote higher uncertainty regions that agents are incentivized to explore. (Right) AirSim-based 3D deployment, where three UAVs cooperatively track multiple moving animal targets modeled with distinct 3D meshes.
Earlier research illuminates separate facets of this challenge but stops short of a unified solution. STAMP leverages temporal attention to trade exploration for revisit frequency but assumes a single robot [3]. MA-G-PPO extends attention mechanisms to multi-agent reinforcement learning (MARL) yet depends on continuous communication and a coarse global occupancy grid [4]. GALOPP accounts for sensing, localisation, and connectivity constraints in heterogeneous teams, though it lacks a principled way of incorporating historical observations when choosing future actions [5]. In our work, we address tasks characterized by highly dynamic spatio-temporal dependencies, where both target motions and agent interactions evolve continuously over time. Inspired by recent findings in spatio-temporal reasoning for dynamic robotics and perception tasks [6, 7], we design a unified transformer architecture that explicitly encodes temporal evolution and spatial relations in a theoretically grounded manner. By jointly modelling uncertainty with GPs and coupling spatial and temporal reasoning inside one transformer, COMPASS closes this gap and produces decentralised policies that adapt online to moving targets.
@@ III Problem Setup
We consider the task of multi-agent persistent monitoring, where a team of $M$ autonomous agents monitors $N$ mobile targets $\mathcal{T}=\{T_{1},...,T_{N}\}$ within a bounded 2D space $\mathcal{W}\subseteq\mathbb{R}^{2}$. Each target $j$ has an evolving state (e.g. location, field distribution) $x_{j}(t)\in\mathcal{X}_{j}$, governed by unknown or stochastic dynamics. Here, $X_{j}\subseteq\mathbb{R}^{2}$ denotes the spatial state space of target $j$, representing all possible positions within the monitoring region. The objective is to maintain accurate belief estimates over all target states throughout the mission.
@@ III-A Belief Modeling with Gaussian Processes
In our implementation, we employ the Matérn kernel [25] with separate spatial and temporal length scales. Sensor observations, represented as binary detections indicating the presence or absence of each target within the sensor’s field of view, are incorporated as noise-free data points $((p_{\text{obs}},t_{\text{obs}}),y_{\text{obs}})$ to incrementally update the GPs.
@@ III-C Sensing and Constraints
Figure 2: Overview of the COMPASS spatio-temporal attention network. Agent observations are first integrated by Gaussian Processes (GPs) to update belief maps. Encoded historical node features are processed by a Temporal Encoder and a Spatial Encoder with positional and masking mechanisms. The fused representation is used by shared Actor and Critic heads for decentralized policy execution, enabling cooperative uncertainty-aware decision-making among agents.
@@ IV-A Problem Modeling and Belief Estimation
At each decision step, agents rely on a belief representation constructed using Gaussian Processes (GPs), as described in Section III-A. These GPs provide probabilistic estimates of each target’s presence across space and time, and we evaluate their posterior mean and variance at every node for both the current time $t$ and a short-term future horizon $t+\delta$. These predictions are concatenated into node-level feature vectors that represent the current and anticipated information landscape. To reflect the deployment state, a binary node presence indicator is appended to each node feature vector, indicating whether it is currently visited by any agent. This representation supports decision-making under uncertainty by informing agents of each other’s locations, enabling coordinated spatio-temporal task allocation via the attention-based policy network. Agents share compact measurement updates during execution: after each measurement cycle, the raw measurements from all agents are aggregated and broadcast globally. Each agent then uses this identical set of observations to independently update its local GP, ensuring the belief maps across the team remain synchronized. Thus, although the policy is executed locally, coordination occurs via this global sharing of raw sensor data.
@@ IV-C Network Architecture
Temporal Encoder
For each node $v$, we maintain a pooled history $\mathcal{H}_{v}=\{\mathbf{h}_{v,t-H+1},\dots,\mathbf{h}_{v,t}\}$. We treat the most recent vector $\mathbf{h}_{v,t}$ as the query, and the full history $\mathcal{H}_{v}$ as keys and values, $\mathbf{W}^{Q},\mathbf{W}^{K},\mathbf{W}^{V}$ are learnable weight matrices, i.e.
$\mathbf{Q}=\mathbf{h}_{v,t}\mathbf{W}^{Q},\quad\mathbf{K}=\mathcal{H}_{v}\mathbf{W}^{K},\quad\mathbf{V}=\mathcal{H}_{v}\mathbf{W}^{V}.$
Multi-head attention then computes a fused embedding
$\begin{split}\mathbf{e}_{\text{temp}}(v,t)&=\mathrm{Concat}(\mathrm{head}_{1},\dots,\mathrm{head}_{N_{h}})\,\mathbf{W}^{O},\\ \mathrm{head}_{i}&=\mathrm{Softmax}\!\bigl(\tfrac{\mathbf{Q}_{i}\mathbf{K}_{i}^{\top}}{\sqrt{d_{k}}}\bigr)\,\mathbf{V}_{i}\end{split}$
This temporally-contextual embedding is then forwarded to the spatial encoder.
@@
To capture spatial correlations, the temporal embeddings are processed by a Transformer encoder that attends across all nodes. Each node representation is augmented with spatial priors, including a Laplacian positional encoding and a current presence embedding. The resulting spatial embedding is then concatenated with the shortest path distance to the nearest agent and projected back to the embedding dimension, yielding the final node feature $e_{\text{final}}(v,t)$.
@@
For each agent, the policy head queries the features of its $K$ neighboring nodes using a single-head attention mechanism. The resulting attention scores form the policy output $\pi(a|s)$, a probability distribution over the candidate action set. In parallel, a decentralized critic head estimates the state value $V(s)$ from the agent’s current node embedding. Both heads are driven by the same shared backbone, enabling joint optimization and decentralized execution.
@@ IV-D Training Settings
We train the shared network using Proximal Policy Optimization (PPO) via centralized training. For data collection, we run $N_{\text{env}}=16$ parallel environments, gathering trajectories of length $T=100$ from all agents into a shared experience buffer for joint gradient updates.
The policy is optimized using PPO’s clipped surrogate objective with a clipping factor of $\epsilon=0.2$:
@@
We estimate advantages using GAE ($\gamma=0.99$, $\lambda=0.95$) and supplement the loss with a value function term and an entropy bonus. The policy is updated using the Adam optimizer with a learning rate of $10^{-4}$, decayed exponentially. Total training took approximately 20 GPU-hours on two NVIDIA A100 GPUs.
@@ V-A Experimental Setup
We conduct all experiments in simulated environments. Evaluations including those in Table I and Table II, are performed in a scripted simulator with visualization. We also validate our framework in AirSim for high-fidelity simulation, as shown in Fig. 3, which captures realistic spatial dynamics and multi-agent coordination in 3D settings. The workspace $\mathcal{W}$ is a continuous 2D area discretized into a graph $\mathcal{G}$, where $K$ nodes are uniformly sampled in the unit square $0,1]^{2}$. Each node is connected to its $k=10$ nearest neighbors to ensure connectivity. Unless otherwise specified, we set $K=200$, $r_{\mathrm{sense}}=0.1$ and use a default speed factor of $0.6$, indicating that targets move at $60\%$ of the agents’ speed.
@@ V-B Performance Analysis and Scalability
Figure 3: Average uncertainty over mission time. COMPASS achieves the fastest and most stable uncertainty reduction compared to baseline methods ($K=200,M=3,N=8$). Solid lines show means; shaded areas denote standard deviations over 20 runs.